find the canonical form of the quadratic form 2x12 +x22 + x32 +2x1x2 – 2x1x3 – 4x2x3

1
Expert's answer
2022-01-20T08:39:47-0500

"2x_1^2+2x_1x_2-2x_1x_3+x_2^2-4x_2x_3+x_3^2=0\\
a_{11}=1,a_{12}=-2,a_{13}=-1,a_{14}=0,a_{22}=1,\\a_{23}=1,a_{24}=0,a_{33}=2,a_{34}=0,a_{44}=0\\
\text{The variant of the equation when converting coordinates are determinants }\\
I_1=a_{11}+a_{22}+a_{33}\\
I_2=\begin{vmatrix}
 a_{11} & a_{12} \\
 a_{12} & a_{22}
\end{vmatrix}+\begin{vmatrix}
 a_{22} & a_{23} \\
 a_{23} & a_{33}
\end{vmatrix}+\begin{vmatrix}
 a_{11} & a_{13} \\
 a_{13} & a_{33}
\end{vmatrix}\\
I_3=\begin{vmatrix}
 a_{11} & a_{12} & a_{13} \\
 a_{12} & a_{22} & a_{23} \\
 a_{13} & a_{23} & a_{33}
\end{vmatrix}\\
I(\lambda)=\begin{vmatrix}
 a_{11}-\lambda & a_{12} & a_{13} \\
 a_{12} & a_{22}-\lambda & a_{23} \\
 a_{13} & a_{23} & a_{33}-\lambda
\end{vmatrix}\\
k_2=\begin{vmatrix}
 a_{11} & a_{14} \\
 a_{14} & a_{44}
\end{vmatrix}+\begin{vmatrix}
 a_{12} & a_{24} \\
 a_{24} & a_{44}
\end{vmatrix}+\begin{vmatrix}
 a_{33} & a_{34} \\
 a_{34} & a_{44}
\end{vmatrix}\\
k_3=\begin{vmatrix}
 a_{11} & a_{12} & a_{14} \\
 a_{12} & a_{22} & a_{24} \\
 a_{14} & a_{24} & a_{44}
\end{vmatrix}+\begin{vmatrix}
 a_{22} & a_{23} & a_{24} \\
 a_{13} & a_{33} & a_{34} \\
 a_{14} & a_{34} & a_{44}
\end{vmatrix}+\begin{vmatrix}
 a_{11} & a_{13} & a_{14} \\
 a_{13} & a_{33} & a_{34} \\
 a_{14} & a_{34} & a_{44}
\end{vmatrix}\\~We~ have\\
I_1=4\\
I_2=\begin{vmatrix}
 1 & -2 \\
 -2 & 1
\end{vmatrix}+\begin{vmatrix}
 1 & 1 \\
 1 & 2
\end{vmatrix}+\begin{vmatrix}
 1 & -1 \\
 -1 & 2
\end{vmatrix}=-1\\
I_3=\begin{vmatrix}
 1 & -2 & -1 \\
 -2 &1 & 1 \\
 -1 & 1 & 2
\end{vmatrix}=-4\\
I(\lambda)=\begin{vmatrix}
 1-\lambda & -2 & -1 \\
 -2 &1-\lambda & 1 \\
 -1 & 1 & 2-\lambda
\end{vmatrix}\\
I(\lambda)=-\lambda^3+4\lambda^2+\lambda-4\\
k_2=0,k_3=0\\
\lambda^3-4\lambda^2-\lambda+4=0\\
by~solving~the~polynomial,~we~get\\
\lambda_1=4 ,\lambda_2=1, \lambda_3=-1\\
the~canonical~form~is\\
x_1^2-x_2^2-4x_3^2=0"